OpenStudy (anonymous):

You are playing a solitaire game in which you are dealt three cards without replacement from a simplified deck of 10 cards (marked 1 through 10) . You win if one of your cards is a 10 or if all of your cards are odd. How many winning cards are there if different orders are different hands? what is your chance of winning?

OpenStudy (anonymous):

5*4*3 + 9*8

OpenStudy (anonymous):

\[P(win) = \frac{132}{720}\]

OpenStudy (anonymous):

how did you get the 9*8 term?

OpenStudy (anonymous):

10 is a must - then 9 free-choices and that leaves 8 free choices

OpenStudy (anonymous):

Soo WHAT ?

OpenStudy (anonymous):

OK you are new - the custom here on this site is , that when problem is solved the asker clicks the blue button "best response" to the one who solved it

OpenStudy (anonymous):

So, I'm still trying to understand the "must" part, and it wasn't my original question... I like your answer, just trying to understand it.

OpenStudy (anonymous):

\[ \color{fuchsia}{\text{if one of your cards is a 10 }} \]

OpenStudy (anonymous):

you start there because that's the definition of a win?

OpenStudy (anonymous):

One of the two definitions

OpenStudy (anonymous):

Btw @aromaboss what about a medal ?

OpenStudy (anonymous):

so, with probability of win = # of winning possibilities vs. total possible hands, you say "if I have won, I must have a 10, and given that I have a 10, I have 9*8 ways to choose my remaining 2 cards since they don't matter", so for the "at least one 10" winner, there are 72 possible winning hands, and 720 total hands? Thanks... not sure why I wasn't getting that at first. I'll give you a medal just for chatting... bad form for aromaboss though... :(

OpenStudy (anonymous):

@JakeV8 human morality is in fact an instrument of fitting into society of peers. See !

OpenStudy (anonymous):

much better :)

OpenStudy (anonymous):

so chances of winning is 132/170 ? any explanation how you reached this answer? whats the answer for "how many winning hands are there if different orders are different hands?

OpenStudy (anonymous):

Semper Fidelis

OpenStudy (anonymous):

am not sure how to give or click medal..guide me.. am new here

OpenStudy (anonymous):

5*4*3 + 9*8 = 5 first odd*4second odd*3third odd + OR (9 choices when 10 you have)*(8 choices for the second NON-ten)

OpenStudy (anonymous):

720 = 10 first card*9 second card*8third card

OpenStudy (anonymous):

@aromaboss... you might have already done it... if not, click the "best response" thing beside on of @Mikael's responses.

OpenStudy (anonymous):

thanks mike i will click

OpenStudy (anonymous):

thank you jake

OpenStudy (anonymous):

CHANCES OF WINNING ANSWER AM NOT CLEAR HOW MANY WINNING HANDS ARE THERE IF DIFFERENT ORDER ARE DIFFERENT HANDS. PLEASE HELP ME.

OpenStudy (anonymous):

I think you already have this, but maybe you aren't seeing it with all the discussion here...

OpenStudy (anonymous):

Total chance of winning = [# of ways to win with odds + # of ways to win with a "10"] all divided by total number of possible hands

OpenStudy (anonymous):

I HAVE NO CLUE ABOUT CARD GAMES..IT IS CONFUSING ME SO MUCH TO INTERPRET.

OpenStudy (anonymous):

as Mikael showed, the # of ways to win with odds are 5*4*3 = 60 and the ways to win with a 10 are 9*8 = 72 So there are 60 + 72 = 132 winning hands And there are 10*9*8 = 720 possible hands So chance of winning is 132 / 720

OpenStudy (anonymous):

The way you get the 5*4*3 is this: How many odds in a 10 card deck where the cards are just numbered "1, 2, 3, etc" Answer: 5 at first... but then you still draw 2 more cards. So 5 odds are possible for the first card, and for each of those, there are 4 remaining odds in the mini-deck. On the third card, there are only 3 remaining odds left. (say for example you drew a 7 first, then a 3, then a 1)

OpenStudy (anonymous):

5 odds x 4 odds x 3 odds = 60 hands that are all odds

OpenStudy (anonymous):

JAKEV8! I GOT THE FEEDBACK FROM MY ANSWER FOR WINNING HANDS 132 AS WRONG...NEED TO EVALUATE WHY?

OpenStudy (anonymous):

HERE WINNING IS...ONE OF CARD IS 10 OR IF ALL OF CARDS ARE ODD... DO NOT UNDERSTAND THIS TWIST

OpenStudy (anonymous):

PLEASE SOME ONE HELP ME. AM CONFUSED

OpenStudy (anonymous):

3/10+5/10*4/9*3/8

OpenStudy (anonymous):

276/720

OpenStudy (anonymous):

Here is the correct Formula for this problem..Hope this helps 5*4*3=60 10=1*9*8+9*1*8+9*8*1=3*9*8=216 10=0 Winning Hands=276 # of Hands=10*9*8=720 Probability of winning=276/720

OpenStudy (anonymous):

this answer proves correct